# FlatQuant: Flatness Matters for LLM Quantization

### v2

@@ B.1 Matrix Inversion and Training Cost
We further present the calibration time required by FlatQuant for the LLaMA family in Table  8. Compared to SpinQuant (Liu et al., 2024b) and QAT methods, FlatQuant requires significantly fewer computational resources and less training time, while delivering superior performance. For weight-only quantization, only transformations related to the linear weights are introduced, resulting in a shorter calibration time compared to weight-activation quantization. Moreover, as discussed in Section 4.2, FlatQuant does not need to be combined with GPTQ to achieve optimal performance, further reducing the calibration overhead.
@@
B.2 More Discussions on Kernel Fusion
@@ Appendix C Additional Experiments
C.1 FlatQuant Leads to Flatness
Figure 8: Flatness and mean squared quantization error (MSE) of different Transformer blocks in LLaMA-3-8B during FlatQuant’s training process. Flatness is calculated as the total flatness of all weights and activations within a Transformer block.
Here we show that with the training objective of Equation 4, the learned affine transformations indeed flatten the distribution of weights and activations, as visualized in Figure 1 and Figure 2.
Quantifying the Flatness.
We quantitatively evaluate the flatness of weights and activations by analyzing their channel magnitude distributions. Specifically, each distribution is represented as a one-dimensional vector $\mathbf{d}$, as depicted in Figure 1. To measure flatness, we compute the mean squared error (MSE) between the observed distribution $\mathbf{d}$ and an idealized perfectly flat distribution $\mathbf{d}^{\prime}$. The flat distribution $\mathbf{d}^{\prime}$ is defined such that all channels possess equal magnitudes and the same $\ell_{2}$ norm as $\mathbf{d}$, i.e., $\mathbf{d}^{\prime}=\frac{\|\mathbf{d}\|_{2}}{\sqrt{N}}\cdot\mathbf{1}_{N}$, where $N$ is the number of channels and $\mathbf{1}_{N}$ is an $N$-dimensional vector with all entries equal to one. The MSE between $\mathbf{d}$ and $\mathbf{d}^{\prime}$ serves as a metric for flatness, with lower MSE values indicating a distribution that is closer to uniform across all channels.
With the above definition, we find that flatness well explains the optimization process of FlatQuant. In Figure 8, we plot the flatness and the training objective in Equation 4 (i.e., the MSE) of different Transformer blocks of LLaMA-3-8B during the training process. It can be observed that as the training loss decreases, the flatness also gets improved. This indicates that FlatQuantlearns better transformation to obtain a flatter distribution which eventually contributes to smaller quantization error, i.e. flatness matters for LLM quantization.
C.2 More Experimental Results
@@
Besides the results of MT-Bench, we present the PPL results and the performance on QA tasks for LLaMA-3.1-8B-Instruct model in Table 10.
@@
To further validate the generality of FlatQuant, we conduct experiments on the Qwen-2.5-Instruct models. FlatQuantachieves near-lossless quantization (e.g., only 0.21% accuracy loss on QA tasks for Qwen-2.5-Instruct-32B). The results on language modeling and QA benchmarks are summarized in Table 11.
@@
KV Cache Quantization.
@@
As introduced in Section 4.4, while our primary focus is on weight-activation quantization, FlatQuant serves as a general framework applicable to various quantization tasks. To further evaluate its versatility, we apply FlatQuant to KV cache only quantization. In this setting, we retain high precision for the rest of the model (including weights and activations) and apply the group-wise asymmetric quantization (with a group size of 128) to keys and values. Table 12 presents the results of KV cache quantization using various bit-widths on the LLaMA-3-8B model. Consistent with previous studies (Hooper et al., 2024; Liu et al., 2024c; Ashkboos et al., 2024), we observe that keys are more sensitive to quantization than values. Furthermore, Table 13 compares FlatQuant with QuaRot for KV cache quantization on LLaMA-2-7B and LLaMA-2-13B models. As shown, FlatQuant delivers superior performance in most cases, particularly for lower-bit (2-3 bits). When both keys and values are quantized to 2 bits, FlatQuant outperforms QuaRot by 2.57 in perplexity for the 7B model.
@@
C.3 Additional Discussions
@@
Unlike weight clipping, which has been widely utilized in LLM quantization, activation clipping has been less explored. Although previous studies (Ashkboos et al., 2024; Liu et al., 2024b) show that activation clipping offers only modest benefits for quantization, our method demonstrates that LCT provides significant improvements. As shown in Table 16, applying our transformations prior to clipping allows for a greater proportion of values to be clipped, resulting in better performance. In contrast, applying LCT before the transformation, similar to the approach used in RTN quantization, yields only limited improvements. This is consistent with prior findings (Dettmers et al., 2022) and is largely due to the presence of severe outliers in activation. We also report results using a QuaRot-style clipping method, with 0.9 as the activation clipping threshold and 0.95 for the KV cache clipping threshold. In summary, the integration of transformations enhances the effectiveness of clipping, indicating that their combination significantly improves weight-activation quantization.
@@
C.4 Experiment Details of Figure 6
@@
C.5 Additional Analyses of Inference Latency
Figure 9: Prefill speedup of LLaMA-2-7B on a sequence length of 2048.
@@ C.5 Additional Analyses of Inference Latency
Total FLOPs of Online Transformations.
Memory Consumption of Online Transformations.
We compute the parameter count of each online transformation below: (1) $\mathbf{P}_{a}$: $2(\sqrt{h_{d}})^{2}$; (2) $\mathbf{P}_{o}$: $a^{2}$; (3) $\mathbf{P}_{h}$: $(h_{d}/a)^{2}$; (4) $\mathbf{P}_{ug}$: $2(\sqrt{h_{d}})^{2}$; (5) $\mathbf{P}_{d}$: $2(\sqrt{h_{i}})^{2}$. The total parameter count in one Transformer block is $4h_{d}+2h_{i}+a^{2}+(h_{d}/a)^{2}$. The additional memory consumption during inference is $2(4h_{d}+2h_{i}+a^{2}+(h_{d}/a)^{2})$ bytes, which only consumes about 0.11MB extra memory space for LLaMA-2-7B.
@@ Appendix E Limitations
Figure 18: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-2-7B. Figure 18(a)-18(c) plot the MSE surface of each method, while Figure 18(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 19: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-2-13B. Figure 19(a)-19(c) plot the MSE surface of each method, while Figure 19(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 20: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-2-70B. Figure 20(a)-20(c) plot the MSE surface of each method, while Figure 20(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 21: The mean squared error (MSE) of quantization across Transformer layers and input sequence in LLaMA-3-70B. Figure 21(a)-21(c) plot the MSE surface of each method, while Figure 21(d) overlays these surfaces by dividing each MSE with that of FlatQuant.

### v3

@@ B.1 Matrix Inversion and Training Cost
We further present the calibration time required by FlatQuant for the LLaMA family in Table  5. Compared to SpinQuant (Liu et al., 2024c) and QAT methods, FlatQuant requires significantly fewer computational resources and less training time, while delivering superior performance. For weight-only quantization, only transformations related to the linear weights are introduced, resulting in a shorter calibration time compared to weight-activation quantization. Moreover, as discussed in Section 4.2, FlatQuant does not need to be combined with GPTQ to achieve optimal performance, further reducing the calibration overhead.
@@
B.2 Overhead Analysis of Affine Transformations
Total FLOPs of Online Transformations.
Memory Consumption of Online Transformations.
We compute the parameter count of each online transformation below: (1) $\mathbf{P}_{a}$: $2(\sqrt{h_{d}})^{2}$; (2) $\mathbf{P}_{o}$: $a^{2}$; (3) $\mathbf{P}_{h}$: $(h_{d}/a)^{2}$; (4) $\mathbf{P}_{ug}$: $2(\sqrt{h_{d}})^{2}$; (5) $\mathbf{P}_{d}$: $2(\sqrt{h_{i}})^{2}$. The total parameter count in one Transformer block is $4h_{d}+2h_{i}+a^{2}+(h_{d}/a)^{2}$. The additional memory consumption during inference is $2(4h_{d}+2h_{i}+a^{2}+(h_{d}/a)^{2})$ bytes, which only consumes about 3.41MB extra memory space for LLaMA-2-7B.
B.3 Detailed Design of Kernel Fusion
@@ Appendix C Additional Experiments
C.1 Results on Other LLM Architectures
@@
Aside from the pre-trained LLaMA models, we also investigate the quantization performance of LLaMA-3.1-8B-Instruct, a representative of the instruction-tuned LLM. The perplexity of language modeling and the accuracy of QA tasks for LLaMA-3.1-8B-Instruct are shown in Table 7, where FlatQuant again outperforms QuaRot by a large margin.
@@
In addition to the series of LLaMA models, we also validate FlatQuant on Qwen-2.5-Instruct, including both the 7B and 32B models. The results on language modeling and QA benchmarks are summarized in Table 8. For the 7B model, FlatQuant achieved a slightly lower average performance compared to the FP16 baseline, with an average score of 68.62. For the 32B model, FlatQuantalso demonstrates competitive performance, achieving an average score of 74.89 (e.g., merely 0.21% drop), which is slightly lower than the FP16 baseline but higher than the QuaRot.
@@
Results on DeepSeek-V3-Base and DeepSeek-R1.
We further scale the evaluation of FlatQuant to DeepSeek-V3-Base and DeepSeek-R1, both of which are large-scale Mixture-of-Experts (MoE) models of 671B parameters. Table 9 presents the results under 4-bit weight and activation quantization. It can be found that FlatQuant maintains strong performance across both LLMs, demonstrating its applicability beyond standard dense LLMs.
Model Quantization C-Eval MMLU AIME2024
DeepSeek V3-Base FP8 90.10 87.10 -
FlatQuant-W4A4 89.59 86.32 -
DeepSeek R1 FP8 - - 79.8
FlatQuant-W4A4 - - 73.3
Table 9: Evaluation results of FlatQuant on DeepSeek V3-Base and DeepSeek R1.
C.2 Results on MT-Bench
Aside from language modeling and question answering, we also evaluate FlatQuant on MT-Bench with LLaMA-3.1-8B-Instruct model in Table 10. We use GPT-4o as the evaluator to justify the ability multi-turn conversation. It can be found that while FlatQuant trails behind the FP16 baseline in coding and STEM, it consistently outperforms QuaRot with GPTQ across all categories, narrowing the gap between the quantized model and the FP16 baseline. Notably, for math problems, FlatQuant matches the FP16 baseline’s score, exceeding QuaRot by 1.9 points.
FP16 8.17 8.10 5.05 7.00 6.10 8.67 8.50 8.91 7.60
QuaRot 7.20 6.90 3.90 5.30 4.05 6.70 6.05 7.80 5.99
FlatQuant 7.95 7.35 4.70 7.20 4.80 7.60 7.20 8.70 6.94
C.3 Extension to More Quantization Settings
Weight-only Quantization
While our primary analysis focuses on joint full weight-activation-kvcache quantization schemes (Section 4.2), FlatQuant demonstrates notable versatility across different quantization paradigms. Table 11 presents the results of the weight-only quantization compared to several state-of-the-art baselines in 4 bits and 3 bits. We adopt per-channel symmetric quantization for weights, maintaining consistency with our full quantization scheme’s weight processing methodology. FlatQuant again obtains leading accuracy compared with leading baselines. Specifically, it outperforms RTN, GTPQ, and AWQ, while also slightly surpassing GPTQ with per-group quantization (group size = 128). Meanwhile, FlatQuant performs comparably to QuIP. Additionally, we report results for FlatQuant when combined with GPTQ, where GPTQ utilizes the same calibration data for both closed-form weight updates and training. The findings remain consistent with the results in joint full weight-activation-kvcache quantization schemes.
LLaMA-3-8B WikiText-2 PPL C4 PPL
W4A16 W3A16 W4A16 W3A16
FP16 6.14 9.45
RTN 8.70 2.2E3 14.00 5.6E3
GPTQ 7.00 13.00 11.80 45.90
GPTQ-g128 6.50 8.20 10.40 13.70
AWQ 7.10 12.80 10.10 16.80
QuIP 6.50 7.50 11.10 11.30
FlatQuant-RTN 6.54 7.78 10.17 12.64
FlatQuant-GPTQ 6.48 7.52 10.28 12.91
Table 11: WikiText-2 and C4 perplexity of weight-only quantizationon on LLaMA-3-8B model.
@@ C.3 Extension to More Quantization Settings
KV Cache Quantization.
To further evaluate its versatility, we apply FlatQuant to KV cache only quantization. In this setting, we retain high precision for the rest of the model (including weights and activations) and apply the group-wise asymmetric quantization (with a group size of 128) to keys and values. Table 12 presents the results of KV cache quantization using various bit-widths on the LLaMA-3-8B model. Consistent with previous studies (Hooper et al., 2024; Liu et al., 2024b; Ashkboos et al., 2024), we observe that keys are more sensitive to quantization than values. Furthermore, Table 13 compares FlatQuant with QuaRot for KV cache quantization on LLaMA-2-7B and LLaMA-2-13B models. As shown, FlatQuant delivers superior performance in most cases, particularly for lower-bit (2-3 bits). When both keys and values are quantized to 2 bits, FlatQuant outperforms QuaRot by 2.57 in perplexity for the 7B model.
@@
Train One and Get More.
Remarkably, we demonstrate that the affine transformations learned from weight-activation quantization can be directly applied to other quantization settings, such as weight-only or KV cache quantization, with surprisingly strong performance. The associated results are presented in Table 15. For instance, the results labeled as “W4” are comparable to those in Table 11 that are specifically trained for weight-only quantization. This significantly saves time when applying FlatQuant to different quantization settings, as only one set of transformation matrices is saved.
W4 A4 KV4 WikiText-2 PPL C4 PPL QA Acc
6.14 9.45 73.23
6.56 10.25 72.92
6.49 10.13 72.20
6.23 9.61 73.43
6.98 11.13 71.23
C.4 Detailed Ablation Study
To better disentangle the contributions of affine transformations and clipping thresholds, we present the full ablation results in Table 16, as an extension of the analysis in Table 3. The results highlight the effectiveness of each component in FlatQuant, with learnable transformations (LT) playing a central role. Moreover, when built upon LT, other components further enhance the overall performance of FlatQuant.
LT PS LCT WikiText-2 C4 ARC-C ARC-E HellaSwag LAMBADA PIQA Winogrande Avg.
1266.60 936.41 25.26 28.62 27.04 1.26 51.80 51.93 30.99
NaN NaN 22.70 25.08 25.04 0.00 49.51 49.57 28.65
1149.08 1490.08 22.95 29.29 27.35 0.60 52.99 50.83 30.67
8197.96 4654.07 25.43 25.72 25.96 0.02 50.49 48.86 29.41
8.50 13.51 44.97 71.38 73.17 67.05 76.88 67.48 66.82
7.95 12.74 44.20 71.89 74.21 68.72 77.15 66.30 67.08
7.11 11.47 49.32 76.14 76.30 72.17 78.89 71.51 70.72
6.98 11.13 50.00 75.80 76.80 72.91 79.16 72.69 71.23
C.5 Additional Discussions
@@
While weight clipping has been widely adopted in LLM quantization, activation clipping remains relatively underexplored. Prior works (Ashkboos et al., 2024; Liu et al., 2024c) have shown that activation clipping alone yields limited benefits, primarily due to the presence of severe outliers. In contrast, our method demonstrates that learnable clipping thresholds (LCT), when applied after transformation, yields significant improvements. As shown in Table 18, applying LCT before transformation, similar to the RTN-style approach, yields only marginal gains. This observation is consistent with prior findings (Dettmers et al., 2022), suggesting that early clipping fails to effectively suppress activation outliers. In our method, the affine transformation redistributes outliers across channels, enabling the subsequent LCT step to more effectively clip a larger portion of extreme values. Crucially, the inverse transformation retains the ability to recover the original scale of informative signals, thereby preserving model quality after quantization. For comparison, we also report results using a QuaRot-style clipping strategy (with thresholds of 0.9 for activations and 0.95 for KV cache values). Overall, these results highlight that the integration of learnable transformations and adaptive clipping significantly enhances the performance of weight-activation quantization.
@@
Effect of Mixed-precision Schemes.
To further evaluate the practicality of FlatQuant, we evaluate its performance under a mixed-precision quantization scheme. Specifically, we explore how integrating FlatQuant into a layer-wise heterogeneous bit-width setting can improve accuracy while retaining high inference speedup. As shown in Table 19, we apply W8A8 quantization to the top 5 important Transformer layers and all down-projection layers, based on their relative importance (Kim et al., 2024). This configuration substantially mitigates the degradation observed in uniform low-bit settings. These results indicate that FlatQuant can be effectively combined with mixed-precision strategies, enhancing its applicability in real-world deployments.
WikiText-2 C4 ARC-C ARC-E HellaSwag LAMBADA PIQA Winogrande Avg
FP16 6.14 9.45 53.50 77.57 79.12 75.51 80.74 72.93 73.23
FlatQuant 6.98 11.13 50.00 75.80 76.80 72.91 79.16 72.69 71.23
+ down_proj_8bits 6.73 10.62 50.00 77.78 77.49 73.96 79.54 70.56 71.55
+ Top5_8bits 6.80 10.82 49.23 76.56 76.54 73.51 79.71 73.95 71.58
+ Top5 & down_proj_8bits 6.61 10.43 51.11 77.74 77.47 74.69 79.92 72.14 72.18
Table 19: The effect of mixed-precision quantization with selectively higher-bit layers in FlatQuant.
C.6 Experiment Details of Figure 6
@@
C.7 Analysis of Inference Memory
To further validate the efficiency of FlatQuant, we provide additional results on its inference-time memory consumption. These experiments complement our theoretical analysis in Appendix B.2, and empirically confirm that the online affine transformations introduced in FlatQuant incur negligible memory overhead.
Table 20 reports the peak memory usage during single-token decoding on a single Transformer layer of the LLaMA-2-7B model, under varying KV cache lengths, with batch size set to 1. We compare standard FP16 inference, INT4 quantization, and our method, FlatQuant. As shown, FlatQuant matches the memory efficiency of INT4 quantization across all sequence lengths, achieving a consistent memory reduction of over 3.3$\times$ compared to FP16. These results indicate that FlatQuant preserves the low memory footprint of INT4 quantization, while introducing no additional memory cost from its online processing.
Sequence Length FP16 (GB) INT4 (GB) FlatQuant (GB) Saving Factor
256 0.393 0.110 0.110 3.58$\times$
512 0.399 0.112 0.112 3.56$\times$
1024 0.411 0.118 0.118 3.48$\times$
2048 0.434 0.130 0.130 3.35$\times$
Table 20: Peak memory usage for decoding a single token on one Transformation layer of LLaMA-2-7B model with KV caches of different lengths and batch size 1.
C.8 Additional Analysis of Inference Latency
@@
Figure 9: Prefill speedup of LLaMA-2-7B on a sequence length of 2048.
Speedup across Sequence Lengths.
In addition to the results presented in Figure 4, we further evaluate the runtime speedup of FlatQuant under varying sequence lengths to better characterize its practical efficiency. Table 21 and Table 22 provide detailed prefill and decoding speedups on LLaMA-3-8B, under different input and KV cache lengths. For the prefill stage with batch size 1, FlatQuant achieves consistent speedups across sequence lengths, achieving 2.12× at length 2048 and 1.80× at length 16384, comparable to INT4 and outperforming QuaRot. In the decoding stage with batch size 64, FlatQuant consistently surpasses QuaRot across all KV cache lengths and closely approaches the efficiency of INT4 quantization. These results demonstrate that FlatQuant maintains its low-overhead advantage across a wide range of generation scenarios, including both short and long contexts.
Prefill Length INT4 QuaRot FlatQuant
2048 2.16$\times$ 1.97$\times$ 2.12$\times$
4096 2.06$\times$ 1.90$\times$ 2.04$\times$
8192 1.94$\times$ 1.79$\times$ 1.92$\times$
16384 1.83$\times$ 1.72$\times$ 1.80$\times$
Table 21: Prefill Speedup on LLaMA-3-8B Compared to FP16 for Different Input Sequence Lengths at Batch Size 1.
KV Cache Length INT4 QuaRot FlatQuant
256 1.38$\times$ 1.09$\times$ 1.24$\times$
512 1.62$\times$ 1.38$\times$ 1.56$\times$
1024 1.70$\times$ 1.61$\times$ 1.63$\times$
2048 1.78$\times$ 1.72$\times$ 1.76$\times$
Table 22: Decoding Speedup on LLaMA-3-8B Compared to FP16 for Different KV Cache Lengths at Batch Size 64.
C.9 Comparison with AffineQuant
AffineQuant also proposed learning an equivalent affine transformation to reduce quantization error. As shown in Table 1, AffineQuant and FlatQuant exhibit markedly different performance. To further clarify why FlatQuant outperforms AffineQuant, how their differences affect quantization performance, and what factors matter most in quantization, we provide additional analysis below.
Improved Expressivity.
Although FlatQuant adopts a Kronecker product with two lightweight matrices, its expressivity remains competitive. As shown in Figure 6, the decomposed transformation achieves accuracy comparable to a full-size affine transformation, suggesting that such a structural constraint does not lead to a loss in functional capacity. In contrast, AffineQuant enforces strictly diagonally dominant transformations to ensure invertibility, which may inadvertently reduce expressivity and make the transformation behave similarly to per-channel scaling (see Figure 7 in the AffineQuant paper).
Applicability to All Linear Layers.
FlatQuant offers broader applicability across the model architecture. Thanks to its lightweight Kronecker structure, it can be efficiently applied to all linear layers with negligible overhead. In contrast, AffineQuant directly learns a full-size affine matrices, which are only feasible to apply to output projection layers due to their high computational cost. For other layers, AffineQuant reverts to simpler per-channel scaling. These practical limitations further constrain its overall effectiveness, whereas FlatQuant maintains a unified and expressive transformation across the entire model.
@@ Appendix E Limitations
Figure 18: The MSE of quantization across Transformer layers and input sequence in LLaMA-2-7B. Figure 18(a)-18(c) plot the MSE surface of each method, while Figure 18(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 19: The MSE of quantization across Transformer layers and input sequence in LLaMA-2-13B. Figure 19(a)-19(c) plot the MSE surface of each method, while Figure 19(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 20: The MSE of quantization across Transformer layers and input sequence in LLaMA-2-70B. Figure 20(a)-20(c) plot the MSE surface of each method, while Figure 20(d) overlays these surfaces by dividing each MSE with that of FlatQuant.
Figure 21: The MSE of quantization across Transformer layers and input sequence in LLaMA-3-70B. Figure 21(a)-21(c) plot the MSE surface of each method, while Figure 21(d) overlays these surfaces by dividing each MSE with that of FlatQuant.